# Simultaneous Uniformization and Algebraic Correspondences

### v1

@@ 4.5. Holomorphic embedding of $\mathrm{Teich}(S_{0,n+1})$ into a space of correspondences
$\mathrm{Teich}(S_{0,n+1})\to\mathbb{C}^{n-1}$
$\Gamma\mapsto R_{\Gamma}\hskip 11.38092pt$
@@
where $\psi_{\rho}$ is a quasiconformal homeomorphism of $\widehat{\mathbb{C}}$ that is conformal on ^$\widehat{\mathbb{C}}\setminus\overline{\mathbb{D}}$ (cf. [Mar16, §5.10]). The quasiconformal maps $\psi_{\rho}$ depend holomorphically on the complex coordinates on the Bers slice $\mathcal{B}(\Gamma_{2})$. We denote the standard complex structure on the Riemann sphere by $\mu_{0}$, and set $\mu_{\rho}:=\psi_{\rho}^{*}(\mu_{0})$. By construction, $\mu_{\rho}$ depends holomorphically on $\rho$, and is $\Gamma_{2}$-invariant. Note that the quasiconformal maps $\psi_{\rho}$ also conjugates $A_{\Gamma_{2}}$ to $A_{\Gamma}$.
@@
Let $R_{\Gamma}$ be as in Subsection 4.4 for a marked group $\Gamma\in\mathrm{Teich}(\Gamma_{2})$. We also denote by $\mathfrak{D}_{\Gamma}$ the Jordan domain satisfying Conditions 12, and 3. We define $\widehat{\mu}_{\Gamma}:=R_{\Gamma_{2}}^{*}(\mu_{\Gamma})$, and note that $\widehat{\mu}_{\Gamma}$ also depends holomorphically on $\Gamma$. The relation $F_{\Gamma_{2}}\circ R_{\Gamma_{2}}\equiv R_{\Gamma_{2}}\circ\eta$ (on $\mathfrak{D}_{\Gamma_{2}}$) and $F_{\Gamma_{2}}$-invariance of $\mu_{\Gamma}$ imply that $\widehat{\mu}_{\Gamma}$ is an $\eta$-invariant Beltrami coefficient. Let $\widehat{\psi}_{\Gamma}$ be a quasiconformal homeomorphism of the sphere solving the Beltrami equation with coefficient $\widehat{\mu}_{\Gamma}$; i.e., $\widehat{\psi}_{\Gamma}^{*}(\mu_{0})=\widehat{\mu}_{\Gamma}$. Then, $\widehat{\psi}_{\Gamma}\circ\eta\circ\widehat{\psi}_{\Gamma}^{-1}$ is a Möbius involution. We normalize $\widehat{\psi}_{\Gamma}$ so that it sends ±$\pm 1,\infty$ to ±$\pm 1,\infty$ (respectively). It then follows that $\widehat{\psi}_{\Gamma}$ depends holomorphically on $\Gamma$, and conjugates $\eta$ to itself. It is now easy to see that $\psi_{\Gamma}\circ R_{\Gamma_{2}}\circ\widehat{\psi}_{\Gamma}^{-1}$ is a quasiregular map of $\widehat{\mathbb{C}}$ preserving the standard complex structure, and hence is a rational map (see Figure 4.5). Further, the rational map $\psi_{\Gamma}\circ R_{\Gamma_{2}}\circ\widehat{\psi}_{\Gamma}^{-1}$ is injective on $\widehat{\psi}_{\Gamma}(\mathfrak{D}_{\Gamma_{2}})$, and we have
@@
The arguments of Proposition 3.2 can also be used to construct conformal matings of higher Bowen-Series maps and factor Bowen-Series maps of the same degree. However, we do not know if such matings admit algebraic descriptions analogous to the one given in Proposition 3.4 (which asserts that matings of two factor Bowen-Series maps is an algebraic function). In other words, we do not know how to lift conformal matings of higher Bowen-Series maps and factor Bowen-Series maps to produce algebraic correspondences uniformizing a pair of genus zero orbifolds.

### v2

@@ 4.5. Holomorphic injection of $\mathrm{Teich}(S_{0,n+1})$ into a space of correspondences
$\boldsymbol{\Psi}:\$$\mathrm{Teich}(S_{0,n+1})\to\mathscr{V}\subset\mathbb{C}^{n-1}$
$\hskip 56.9055pt\Gamma\mapsto R_{\Gamma}$
@@
where $\psi_{\rho}$ is a quasiconformal homeomorphism of $\widehat{\mathbb{C}}$ that is conformal on $\widehat{\mathbb{C}}\setminus\overline{\mathbb{D}}$ (cf. [Mar16, §5.10]). The quasiconformal maps $\psi_{\rho}$ depend holomorphically on the complex coordinates on the Bers slice $\mathcal{B}(\Gamma_{2})$. We denote the standard complex structure on the Riemann sphere by $\mu_{0}$, and set $\mu_{\rho}:=\psi_{\rho}^{*}(\mu_{0})$. By construction, $\mu_{\rho}$ depends holomorphically on $\rho$, and is $\Gamma_{2}$-invariant. Note that the quasiconformal maps $\psi_{\rho}$ also conjugates $A_{\Gamma_{2}}$ to $A_{\Gamma}$.
@@
Let $R_{\Gamma}$ be as in Subsection 4.4 for a marked group $\Gamma\in\mathrm{Teich}(\Gamma_{2})$. We also denote by $\mathfrak{D}_{\Gamma}$ the Jordan domain satisfying Conditions 12, and 3. We define $\widehat{\mu}_{\Gamma}:=R_{\Gamma_{2}}^{*}(\mu_{\Gamma})$, and note that $\widehat{\mu}_{\Gamma}$ also depends holomorphically on $\Gamma$. The relation $F_{\Gamma_{2}}\circ R_{\Gamma_{2}}\equiv R_{\Gamma_{2}}\circ\eta$ (on $\mathfrak{D}_{\Gamma_{2}}$) and $F_{\Gamma_{2}}$-invariance of $\mu_{\Gamma}$ imply that $\widehat{\mu}_{\Gamma}$ is an $\eta$-invariant Beltrami coefficient. Let $\widehat{\psi}_{\Gamma}$ be a quasiconformal homeomorphism of the sphere solving the Beltrami equation with coefficient $\widehat{\mu}_{\Gamma}$; i.e., $\widehat{\psi}_{\Gamma}^{*}(\mu_{0})=\widehat{\mu}_{\Gamma}$. Then, $\widehat{\psi}_{\Gamma}\circ\eta\circ\widehat{\psi}_{\Gamma}^{-1}$ is a Möbius involution. We normalize $\widehat{\psi}_{\Gamma}$ so that it sends $\pm 1,\infty$ to $\pm 1,\infty$ (respectively). It then follows that $\widehat{\psi}_{\Gamma}$ depends holomorphically on $\Gamma$, and conjugates $\eta$ to itself. It is now easy to see that $\psi_{\Gamma}\circ R_{\Gamma_{2}}\circ\widehat{\psi}_{\Gamma}^{-1}$ is a quasiregular map of $\widehat{\mathbb{C}}$ preserving the standard complex structure, and hence is a rational map (see Figure 4.5). Further, the rational map $\psi_{\Gamma}\circ R_{\Gamma_{2}}\circ\widehat{\psi}_{\Gamma}^{-1}$ is injective on $\widehat{\psi}_{\Gamma}(\mathfrak{D}_{\Gamma_{2}})$, and we have
@@
The group $\Gamma_{2}$ merely plays the role of a base-point in the proof of Proposition 4.3; while it is used to put a complex structure on $\mathrm{Teich}(S_{0,n+1})$, the image of the map $\boldsymbol{\Psi}$ does not depend on $\Gamma_{2}$. In fact, for any $\Gamma_{2}^{\prime}\in\mathrm{Teich}(S_{0,n+1})$, the corresponding Bers slices $\mathcal{B}(\Gamma_{2})$ and $\mathcal{B}(\Gamma_{2}^{\prime})$ are naturally biholormorphic, and hence one can use any such Bers slice $\mathcal{B}(\Gamma_{2}^{\prime})$ to equip $\mathrm{Teich}(S_{0,n+1})$ with a complex structure.
Combining Propositions 4.2 and 4.3, we readily obtain the following:
Theorem 4.4.
Let $n\geq 2$. There exists an injective holomorphic map
$\boldsymbol{\Psi}:\$$\mathrm{Teich}(S_{0,n+1})\to\mathscr{V}\subset\mathbb{C}^{n-1}$
$\hskip 56.9055pt\Gamma\mapsto R_{\Gamma},$
where $\mathscr{V}$ is the (n2)(n-2)-dimensional algebraic variety defined by Equation (4.2), such that the correspondence $\mathfrak{C}$ defined by Equation (4.3) simultaneously uniformizes the (marked) $(n+1)$-times punctured sphere $\mathbb{D}/\Gamma$ and the Hecke orbifold (having one puncture, an order 2 orbifold point and an order $2n$ orbifold point).
We conclude with several questions and comments.
Non-singularity and holomorphic embedding of $\mathrm{Teich}(S_{0,n+1})$ into a space of correspondences
If $\boldsymbol{\Psi}(\mathrm{Teich}(S_{0,n+1}))$ contains no singular point of $\mathscr{V}$, then it is a complex manifold of dimension n2n-2, and hence the injective holomorphic map $\boldsymbol{\Psi}$ would be a biholomorphism onto its image. To this end, we ask:
Question 4.5.
Is the variety $\mathscr{V}$ defined by Equation (4.2) non-singular? If it is not, does the image of $\boldsymbol{\Psi}$ contains a singular point of $\mathscr{V}$?
Boundedness of Bers slices
Classical Bers slices are pre-compact in the $\operatorname{PSL}_{2}(\mathbb{C})$-character variety [Ber70]. As mentioned the introduction, holomorphic embeddings of Bers slices of genus zero orbifolds in spaces of algebraic correspondences was constructed in [MM23c], with the property that the correspondences in the image are matings of the corresponding surfaces and the polynomial $z^{d}$. Pre-compactness of these images was proved in [LMM25]. It is natural to ask the same question in the current setting.
Question 4.6.
1) Is the image of the map $\boldsymbol{\Psi}$ bounded in the variety $\mathscr{V}$?
2) Analyze the dynamics of the correspondences lying on the boundary of the image of $\boldsymbol{\Psi}$.
Furthermore, many of the questions posed in [LMM25, Question 1.2] make sense in the present setting too.
Higher Bowen-Series maps, and algebraic correspondences
The arguments of Proposition 3.2 can also be used to construct conformal matings of higher Bowen-Series maps and factor Bowen-Series maps of the same degree. However, we do not know if such matings admit algebraic descriptions analogous to the one given in Proposition 3.4 (which asserts that the mating of two factor Bowen-Series maps is an algebraic function). We raise the following question:
Question 4.7.
Can the conformal matings of higher Bowen-Series maps and factor Bowen-Series maps (of equal degree) be lifted to produce algebraic correspondences uniformizing a pair of genus zero orbifolds?
Quasi-Fuchsian locus, and boundary correspondences
In Theorem 1.1, we consrtucted correspondences that combine genus zero $\Sigma_{1},\Sigma_{2}$ satisfying certain compatibility conditions. While we focussed on the special case where $\Sigma_{1}$ is the Hecke orbifold in Section 4, one can consider the situation where both $\Sigma_{1}$ and $\Sigma_{2}$ have non-trivial Teichmüller spaces. In this general setup, the construction of this section would give rise to a holomorphic injection from $\mathrm{Teich}(\Sigma_{1})\times\mathrm{Teich}(\Sigma_{2})$ into an appropriate space of algebraic correspondences that combine marked surfaces in $\mathrm{Teich}(\Sigma_{1})$, $\mathrm{Teich}(\Sigma_{2})$. The following problem is of considerable interest.
Question 4.8.
1) Prove an analog of Thurston’s Double Limit Theorem for correspondences lying in the image of $\mathrm{Teich}(\Sigma_{1})\times\mathrm{Teich}(\Sigma_{2})$ under the above holomorphic injection.
2) Are there correspondences on the boundary of the image of $\mathrm{Teich}(\Sigma_{1})\times\mathrm{Teich}(\Sigma_{2})$ (under the above holomorphic injection) that are matings of two singly degenerate groups lying on $\partial\mathrm{Teich}(\Sigma_{1})$, $\partial\mathrm{Teich}(\Sigma_{2})$?